OpenStudy (anonymous):

I need help solving an equation using the "u" |2-2t|+1=7

OpenStudy (anonymous):

I'm not sure what you mean by using the "u" but solving this I would do this: |2-2t|+1=7 |2-2t|=6 2-2t=6 and 2-2t=-6 -2t=4 and -2t=-8 t = -2 and t = 4

OpenStudy (anonymous):

the U is "union" and is translated into english as "or"

OpenStudy (anonymous):

TBates use of the word "and" is mathematically an error; we might say that in common english, but in math-speak -2 union 4 is -2 or 4.

OpenStudy (anonymous):

the algebraic work above is correct, but absolute value equations are translated using the union or U.

OpenStudy (anonymous):

Ah, that makes sense. Most of the time I write x = -2, 4 and the comma is generally regarded as being "or" anyway but I suppose "mathematically grammatically" I should use union of sets to give the answer.

OpenStudy (anonymous):

think of it this way: what is the intersection of the sets {-2} and {4}? answer= null set compared to the union of {-2} U {4}? answer= {-2, 4}

OpenStudy (anonymous):

liljenredd, you will use U or union or "or" anytime you have an absolute value equation, or absolute value expression > positive number; when you have abs. values expression < positive number you will use "and" or intersection

OpenStudy (anonymous):

the U sub

OpenStudy (anonymous):

i don't understand you roast post "sub"?

OpenStudy (anonymous):

no roast but last!!!!!!

OpenStudy (anonymous):

I understand I was just telling Tbates what I was trying to type

OpenStudy (anonymous):

do you want an example?

OpenStudy (anonymous):

but you all are right

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

can you show me how you got that answer

OpenStudy (anonymous):

Example: solve |2x-3| > 6 (solution to follow on next post)

OpenStudy (anonymous):

by definition of absolute value -(2x-3)>6 or 2x-3>6 2x-3<-6 or 2x-3>6 2x<-3 or 2x>9 (add 3 to both sides of both inequalities) x<-3/2 or x>9/2 (divide by 2) the graphical solution isthe interval notation is\[(-\infty, -3/2) \cup (9/2, \infty)\]